1) 2x+3x+4x=2x
2) x/3=5x/4

Answer by Alan3354(69443)

1) 2x+3x+4x=2x
3x+4x = 0
x = 0
--------------
2) x/3=5x/4
Cross multiply
4x = 15x
-11x = 0
x = 0